OpenStudy (anonymous):

17-[22÷21-x(y))] if x=2 and y=5

OpenStudy (anonymous):

17-11-xy if x=2 and y =5 17-11-(2)(5) 17-11-10 = ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

no, 17-11-10 = -4

OpenStudy (anonymous):

i got 15 for it

OpenStudy (anonymous):

17-11=15

OpenStudy (anonymous):

depends on where the parentheses are. is it 17-[22÷(21-x(y))] or 17-[22/21 -xy]

OpenStudy (anonymous):

wait my bad

OpenStudy (anonymous):

and no, 17-11 = 6

OpenStudy (anonymous):

it is 17-[22÷(21-x(y))]

OpenStudy (anonymous):

then the answer is 15

OpenStudy (anonymous):

oh, yea. 15 sorry hehe

OpenStudy (anonymous):

haa